OpenStudy (anonymous):

What is 6/15 as a decimal?

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

nice try liana :)

OpenStudy (anonymous):

i knew the answer

OpenStudy (anonymous):

oh hi praty

OpenStudy (anonymous):

6/15 = 0.4

OpenStudy (anonymous):

lol

OpenStudy (pratyush5):

@jinxhead20 Rise and shine :) morning :P

OpenStudy (anonymous):

so btw 6/15 = 0.4

OpenStudy (anonymous):

u copying me i wrote that first

OpenStudy (anonymous):

@liana123??? *facepalm*

OpenStudy (marigirl):

@deadpool123 the best way to do this question without a calc: \[\frac{ 6 } {15 } = \frac{ 2 }{ 5 }\] . this is a simplified fraction since we got that sorted what we can do is multiply the top and bottom by 20 \[\frac{ 2 }{ 5 } \times \frac{ 20 }{ 20 }=\frac{ 40 }{ 100 }\] and we know that is 40% or 0.4

OpenStudy (pratyush5):

6/15 = 0.4 Lets join the party baby :P